# Do Not Design, Learn: A Trainable Scoring Function for Uncertainty Estimation in Generative LLMs

### v1

@@ Abstract
In this work, we introduce the Learnable Response Scoring Function (LARS) for Uncertainty Estimation (UE) in generative Large Language Models (LLMs). Current scoring functions for probability-based UE, such as length-normalized scoring and semantic contribution-based weighting, are designed to solve specific aspects of the problem but exhibit limitations, including the inability to handle biased probabilities and under-performance in low-resource languages like Turkish. To address these issues, we propose LARS, a scoring function that leverages supervised data to capture complex dependencies between tokens and probabilities, thereby producing more reliable and calibrated response scores in computing the uncertainty of generations. Our extensive experiments across multiple datasets show that LARS substantially outperforms existing scoring functions considering various probability-based UE methods.
**footnotetext: Equal Contribution
@@ 1 Introduction
Figure 1: (Left) Answer generation process using a generative LLM. (Mid Left) LARS overview. It utilizes the question, answer tokens, and token probabilities. Token probabilities are fed to LARS model as special probability tokens. (Mid Right) Illustration of few-hot represented embedding vectors of probability tokens. (Right) Summary of probability-based UE methods where they take different sampled answer scores and output a single UE value.
@@
Recent years have seen a transformative shift in AI due to the emergence of generative Large Language Models (LLMs). Their near-human capabilities in understanding, generating, and processing information have revolutionized human-machine interactions and facilitated their integration across various industries such as healthcare, law, finance, and marketing Ye et al. (2023); OpenAI (2023); Touvron et al. (2023); Huang et al. (2023). Given that LLMs can sometimes generate misleading or erroneous outputs, it is crucial to evaluate how much reliance should be placed on their responses. Tools such as hallucination detection Li et al. (2023), fact verification Wang et al. (2024), and Uncertainty Estimation (UE) Malinin and Gales (2021) are essential for assessing the correctness of model responses. The field of Uncertainty Estimation, well-established in classification tasks, has recently been adapted to generative LLMs. Recent studies Kuhn et al. (2023) demonstrate that these adaptations can effectively predict incorrect LLM outputs without the need for external feedback.
UE methods can be broadly categorized into two approaches. Probability-based methods Malinin and Gales (2021); Kuhn et al. (2023) utilize token probabilities externally to predict uncertainty. In contrast, non-probability-based methods Lyu et al. (2024); Chen et al. (2024) employ heuristics that do not rely on token probabilities. This work focuses exclusively on probability-based methods, with a discussion of related works presented in Section 2.
@@
A fundamental challenge in UE of LLMs with probability-based methods is the necessity to aggregate multiple token probabilities into a single score. To this end, existing methods typically employ a scoring function. A common scoring function is Length-Normalized Scoring (LNS), which calculates the mean of log probabilities, as employed by Malinin and Gales (2021); Kuhn et al. (2023), to mitigate bias in longer generations. Subsequent approaches by Bakman et al. (2024); Duan et al. (2024) introduce heuristics that prioritize semantically important tokens by assigning higher weights to them, rather than simply averaging as in LNS. However, these scoring functions, largely heuristic in design, often overlook potential pitfalls. In this work, we critically analyze the weaknesses of the existing scoring functions and introduce Learnable Response Scoring Function (LARS), which learns a scoring function from supervised data.
@@
We summarize our main contributions as follows: 1. We experimentally demonstrate the limitations of existing scoring functions in terms of their calibration and performance in low-resource languages. 2. We introduce a novel off-the-shelf scoring function, LARS, which is learned directly from supervised data. 3. We validate the superiority of LARS over existing baselines across three different datasets and provide an analysis of its components to rationalize the effectiveness of LARS.
2 Background
Uncertainty Estimation (UE).
@@ 2 Background
Uncertainty Estimation (or Quantification) in generative Large Language Models (LLMs) addresses the challenge of predicting a model’s uncertainty regarding a given sequence or question. An effective UE method assigns a lower score (indicating less uncertainty) to questions where the model is likely to provide the correct answer, and a higher score otherwise. Mathematically, we have $UE(\theta,x_{1})<UE(\theta,x_{2})$ if the most probable generation of model $\theta$ for question $x_{1}$ is more likely to be correct than for question $x_{2}$. Previous works formulate this approach for closed-ended questions with well-defined ground truths Malinin and Gales (2021); Kuhn et al. (2023); Bakman et al. (2024); Duan et al. (2024).
Related Works.
UE has recently become a topic of significant interest, leading to the proposal of various methods. These methods can be broadly categorized into four types: 1. Self-checking methods: The model evaluates its own generation correctness using different strategies Kadavath et al. (2022); Manakul et al. (2023); Li et al. (2024); Luo et al. (2023); Zhao et al. (2023). 2. Output consistency methods: Uncertainty is predicted by examining the consistency of various outputs for a given question Lyu et al. (2024); Lin et al. (2023); Zhang et al. (2024); Ulmer et al. (2024); Elaraby et al. (2023). 3. Internal state examination methods: The activations of the model are analyzed to predict the model errors Chen et al. (2024). 4. Token probability-based methods: These methods utilize token probabilities to estimate uncertainty Malinin and Gales (2021); Kuhn et al. (2023); Bakman et al. (2024); Duan et al. (2024). These methods can be used in conformal prediction frameworks, which offer theoretical guarantees for model correctness Deutschmann et al. (2024); Quach et al. (2023); Yadkori et al. (2024). In this work, we focus on improving token probability-based methods by proposing a learnable scoring function.
Token Probability-based Methods.
@@
(Malinin and Gales, 2021) formally proposes using sequence probability as the generation’s probability for a given question $\mathbf{x}$ and a model parameterized by $\theta$. This is mathematically defined as follows:
@@
where $P(\mathbf{s}|\mathbf{x},\theta)$ is the sequence probability for the generated sequence $\mathbf{s}$, and $s_{<l}\triangleq\{s_{1},s_{2},\ldots,s_{l-1}\}$ represents the tokens generated before $s_{l}$. This sequence probability is used in entropy calculation $\mathcal{H}(\mathbf{x},\theta)$ by making a Monte Carlo approximation, which requires multiple answer sampling for the given question:
@@
where $\mathbf{s}_{b}$ is a sampled generation to the question $\mathbf{x}$. Later Kuhn et al. (2023) improves the entropy by utilizing the semantic meaning of the sampled generations. They cluster the generations with the same meaning and calculate entropy using the generation probabilities associated with each cluster:
$SE(\mathbf{x},\theta)=-\frac{1}{|C|}\sum_{i=1}^{|C|}\ln P(\mathrm{c}_{i}|\mathbf{x},\theta),$ (3)
where $\mathrm{c}_{i}$ refers to each semantic cluster and $C$ is the set of all clusters. Notably, Aichberger et al. (2024) enhances semantic entropy by enabling the model to generate semantically more diverse outputs.
Both Malinin and Gales (2021) and Kuhn et al. (2023) observe that sequence probability in (1) is biased against longer generations. To address this, they use length-normalized scoring as follows:
@@
where $w(\mathbf{s},\mathbf{x},L,l)$ is the weight of the $l$th token assigned by MARS or TokenSAR. These scoring functions aim to give more weight to tokens that directly answer the question and are calibrated such that if a generation is likely to be incorrect, they yield a lower score, and vice versa. Our goal in this work is to enhance this calibration by learning the scoring function directly from the data.
@@ 3 Shortcomings of Existing Scoring Functions
In this section, we critically and empirically analyze the shortcomings of existing scoring functions, namely Length-Normalized Scoring (LNS), MARS, and TokenSAR.
@@
Existing scoring functions are designed to address particular challenges within the UE problem domain. For instance, LNS mitigates length bias, whereas MARS and TokenSAR focus on reducing the impact of non-essential token probabilities. However, the complexities of designing an optimal scoring function may not be immediately evident. Typically, scoring functions involve a dot product of log probabilities and assigned weights, but alternative formulations could provide more finely calibrated estimations. Additionally, these existing functions may not adequately capture complex dependencies between tokens, such as grammatical and semantic interactions De Marneffe and Nivre (2019). While MARS attempts to address this by weighting phrases rather than individual tokens, it only partially solves the problem and fails to capture deeper dependencies. Lastly, both MARS and TokenSAR apply normalization on their weights $w(\mathbf{s},\mathbf{x},L,l)$, through methods like sum-normalization (TokenSar) or softmax (MARS). These design choices directly impact the model’s output, potentially making the model converge to sub-optimal points.
Figure 2: Average accuracy and probability assignments of LLama2-7b model to specific entities in TriviaQA.
@@
Existing scoring functions often directly utilize token probabilities, which can exhibit biases against specific types of entities. To explore this issue, we conducted an experiment with Llama2-7b Touvron et al. (2023) using the TriviaQA dataset Joshi et al. (2017). We posed questions from TriviaQA to the model and analyzed the probabilities assigned to tokens in the answer representing different entity types such as person names, organizations, and dates. Additionally, we assessed the accuracy of the model across these categories. As presented in Figure 2, although the model shows comparable accuracy for date and person entities, it assigns higher probabilities to tokens associated with dates. This finding suggests a notable positive bias towards date entities. Similar patterns can be observed in other entities. Such differences in probability assignment highlight the need for recalibration across entities, a feature that current scoring functions fail to adequately address.
Low-Resource Language Challenges.
MARS and TokenSAR are dependent on existing NLP tools for implementation. Specifically, TokenSAR uses a sentence similarity model Duan et al. (2024), and MARS relies on a QA evaluator model Bulian et al. (2022). These models may not be readily available for some low-resource languages. Moreover, the design of MARS and TokenSAR is primarily oriented towards English. This orientation can be challenging when these tools are applied to languages that are morphologically distinct from English, such as Turkish Göksel and Kerslake (2005). In Section 5.4, we experimentally demonstrate that existing methods do not yield comparable improvements in Turkish (compared to English).
@@ 4 LARS: Learnable Response Scoring
Let $f$ denote the scoring function, which accepts three arguments: the input prompt $\mathbf{x}=\{x_{1},x_{2},\ldots,x_{N}\}$, the generated sequence $\mathbf{s}=\{s_{1},s_{2},\ldots,s_{L}\}$, and the corresponding probability vector $\mathbf{p}=\{p_{1},p_{2},\ldots,p_{L}\}$, where $p_{i}$ represents the probability of token $s_{i}$. The function $f$ outputs a real number $o$. In token probability-based methods, it is desirable for $o$ to be lower when the generation $\mathbf{s}$ is more likely to be incorrect, improving the model’s uncertainty estimation. As discussed in Section 3, manually designing an effective scoring function is a challenging endeavor. Thus, we propose making the scoring function $f$ directly learnable through supervised data.
@@
We construct a calibration set to train our scoring function, $f_{w}$, which is parameterized by $w$. This calibration set comprises 4-tuples: input prompt $\mathbf{x}$, generated sequence $\mathbf{s}$, probability vector $\mathbf{p}$, and binary ground truth label $g$. The label $g$ indicates whether $\mathbf{s}$ is a correct response to $\mathbf{x}$. To optimize the parameters of $f_{w}$, we employ the binary cross-entropy loss, denoted by $L$, applied as follows:
$L(f_{w}(\mathbf{x},\mathbf{s},\mathbf{p}),g).$
To train the scoring function $f_{w}$, we start with the pre-trained RoBERTa model Liu et al. (2019) and augment it by adding a linear layer that outputs a single logit. The input format for the LARS model is structured as follows: initial prompt $\mathbf{x}$, followed by a series of response tokens $\mathbf{s}=\{s_{1},s_{2},\ldots,s_{L}\}$. Each response token $s_{i}$ is immediately succeeded by a special probability token $\tilde{p_{i}}$. This probability token $\tilde{p_{i}}$ is associated with the probability $p_{i}$.
The model incorporates a total of $k$ distinct probability tokens, each corresponding to a specific partition of the [0, 1] probability range. These partitions are mutually exclusive, cover the entire probability range, and are determined based on the quantiles of the probabilities in the calibration dataset. The probability token $\tilde{p_{i}}$ for $p_{i}$ is selected according to the partition into which $p_{i}$ falls.
The embedding vectors of probability tokens are structured by few-hot encoding approach. Assuming the pretrained model has an input dimension $d$, $r$-th probability token will be represented by setting the vector positions from $(r-1)\times\frac{d}{k}$ to $r\times\frac{d}{k}$ to 1, while all other positions are set to 0. To ensure consistency with the pretrained model’s token embedding norms, we scale these probability vectors by a fixed divisor. $f_{w}$ is visualized in Figure 1.
With this architecture and input strategy, we enable our scoring function to accurately associate each probability $p_{i}$ with its corresponding token $s_{i}$. By representing $p_{i}$ using a few-hot vector format, the scoring function effectively utilizes probability information in a manner analogous to conditional image generation tasks van den Oord et al. (2016). Additionally, using a pretrained model allows the scoring function to grasp the linguistic dependencies and semantic nuances of the tokens. This capability may be crucial in yielding a well-calibrated scoring function to properly employ the probabilities of certain entities, as discussed in Section 3.

### v2

@@ Abstract
Uncertainty estimation (UE) of generative large language models (LLMs) is crucial for evaluating the reliability of generated sequences. A significant subset of UE methods utilize token probabilities to assess uncertainty, aggregating multiple token probabilities into a single UE score using a scoring function. Existing scoring functions for probability-based UE, such as length-normalized scoring and semantic contribution-based weighting, are designed to solve certain aspects of the problem but exhibit limitations, including the inability to handle biased probabilities and complex semantic dependencies between tokens. To address these issues, in this work, we propose Learnable Response Scoring (LARS) function, a novel scoring function that leverages supervised data to capture complex dependencies between tokens and probabilities, thereby producing more reliable and calibrated response scores in computing the uncertainty of LLM generations. Our comprehensive experiments across question-answering and arithmetical reasoning tasks with various datasets demonstrate that LARS significantly outperforms existing scoring functions, achieving improvements of up to 16% AUROC score.
@@ 1 Introduction
Figure 1: (Left) Answer generation using a generative LLM. (Mid Left) Overview of the proposed scoring function LARS. It utilizes the question, answer tokens, and token probabilities. Token probabilities are fed to LARS model as special probability tokens. (Mid Right) Illustration of few-hot represented embedding vectors of probability tokens. (Right) Overview of probability-based UE methods taking different sampled answer scores such as LNS Malinin and Gales (2021), MARS Bakman et al. (2024), or LARS (this work), and outputting a single UE value.
@@
Recent years have seen a transformative shift in AI with the rise of generative Large Language Models (LLMs). Their near-human capabilities in comprehension, generation, and information processing have revolutionized human-machine interactions, driving widespread adoption across industries such as healthcare, law, finance, and marketing Ye et al. (2023); OpenAI (2023); Touvron et al. (2023); Huang et al. (2023). Given that LLMs can sometimes generate misleading or erroneous outputs, it is crucial to evaluate how much reliance should be placed on their responses. Detecting unreliable, factually incorrect, or irrelevant outputs from LLMs is studied under the topic of hallucination detection Li et al. (2023). Methods such as fact verification Wang et al. (2024); Chern et al. (2023), cross examination Cohen et al. (2023) and Uncertainty Estimation (UE) Malinin and Gales (2021) serve as tools for hallucination detection.
The field of UE, well-established in classification tasks, has recently been adapted to generative LLMs. In the context of generative LLMs, UE is used to assess the model’s reliability for a given query Kuhn et al. (2023). UE methods are particularly valuable as they differ from other hallucination detection approaches by not relying on external resources, such as internet search tools Chern et al. (2023) or a teacher model Cohen et al. (2023). UE methods in generative LLMs can be broadly categorized into two categories: 1) Probability-based methods Malinin and Gales (2021); Kuhn et al. (2023) that utilize token probabilities externally to predict uncertainty. 2) Non-probability-based methods Lyu et al. (2024); Chen et al. (2024) that employ heuristics without relying on token probabilities for estimation. This work focuses on probability-based methods due to their widespread use and promising performance in UE Bakman et al. (2024); Duan et al. (2024); Kuhn et al. (2023), as well as their applicability to closed-source API models where token probabilities are accessible OpenAI (2023).
@@
Probability-based UE in LLMs requires aggregating multiple token probabilities into a single score, which can be done through a scoring function. Length-Normalized Scoring (LNS) Malinin and Gales (2021); Kuhn et al. (2023) is a common approach, which calculates the mean of log probabilities of an LLM’s output to mitigate bias in longer generations. Subsequent approaches by Bakman et al. (2024); Duan et al. (2024) introduce heuristics that prioritize semantically important tokens by assigning higher weights to them, rather than simply averaging as in LNS. However, these scoring functions, largely heuristic in design, often overlook potential pitfalls such as biased probabilities and complex dependencies between tokens. In this work, we critically analyze the weaknesses of the existing scoring functions and introduce a novel scoring function that leverages supervised data to produce more calibrated scores for UE in LLMs.
@@
We summarize our main contributions as follows: (1) We discuss the limitations of existing scoring functions of UE from three different perspectives including biased probabilities, token dependencies, and applicability to other languages rather than English. (2) We introduce a novel off-the-shelf scoring function, Learnable Response Scoring (LARS), which is learned directly from supervised data (visualized in Figure 1). (3) We validate the superiority of LARS over existing baselines across three QA datasets, a mathematical reasoning task, and four different languages. LARS outperforms SOTA scoring functions by up to 16% in AUROC and 45% in PRR. Additionally, we analyze its components to explain the effectiveness of LARS.
2 Preliminaries
@@
Uncertainty Estimation in Generative LLMs addresses the challenge of predicting a model’s uncertainty regarding a given input sequence or question. In the context of closed-ended QA and mathematical reasoning tasks, an effective UE method assigns a lower score (indicating less uncertainty) to questions where the model is likely to provide the correct answer (reliable output), and a higher score otherwise. Mathematically, we have $\mathrm{UE}(\theta,x_{1})<\mathrm{UE}(\theta,x_{2})$ if the most probable generation of model $\theta$ for question $x_{1}$ is more likely to be correct than for question $x_{2}$ Malinin and Gales (2021); Kuhn et al. (2023); Duan et al. (2024).
@@
Token Probability-based Methods use token probabilities to estimate the model uncertainty. This estimation requires aggregating multiple token probabilities into a single score. In their foundational work, Malinin and Gales (2021) formalize the generation’s probability for a given question $\mathbf{x}$ and a model parameterized by $\theta$ using the sequence probability. This is defined as follows:
@@
where $P(\mathbf{s}|\mathbf{x},\theta)$ is the probability for the generated sequence $\mathbf{s}$ (of length $L$), and $s_{<l}\triangleq\{s_{1},s_{2},\ldots,s_{l-1}\}$ represents the tokens generated before token $s_{l}$. This sequence probability is used in entropy calculation $\mathcal{H}(\mathbf{x},\theta)$ by making a Monte Carlo approximation, which requires multiple answer sampling for the given question:
@@
where $\mathbf{s}_{b}$ is a sampled generation to the question $\mathbf{x}$. Later Kuhn et al. (2023) improve the entropy by utilizing the semantics of the sampled generations. They cluster the generations with the same meaning and calculate entropy using the generation probabilities associated with each cluster:
$\mathrm{SE}(\mathbf{x},\theta)=-\frac{1}{|C|}\sum_{i=1}^{|C|}\ln P(\mathrm{c}_{i}|\mathbf{x},\theta),$ (3)
where $\mathrm{c}_{i}$ refers to each semantic cluster and $C$ is the set of all clusters. Notably, Aichberger et al. (2024) enhance semantic entropy by enabling the model to generate semantically more diverse outputs.
Both Malinin and Gales (2021) and Kuhn et al. (2023) observe that sequence probability in (1) is biased against longer generations. To address this, they use a length-normalized scoring as follows:
@@
where $w(\mathbf{s},\mathbf{x},L,l)$ is the weight of the $l$-th token assigned by MARS or TokenSAR. These scoring functions aim to give more weight to tokens that directly answer the question and are calibrated such that if a generation is likely to be incorrect, they yield a lower score, and vice versa. Our goal in this work is to enhance this calibration by learning the scoring function directly from the data.
@@ 3 Shortcomings of Existing Scoring Functions
In this section, we discuss the shortcomings of scoring functions: LNS, MARS, and TokenSAR.
@@
Existing scoring functions are designed to address particular challenges within the UE problem domain. For instance, LNS mitigates length bias, whereas MARS and TokenSAR focus on reducing the impact of non-essential token probabilities. However, the complexities involved in designing an optimal scoring function may not be immediately apparent. Typically, scoring functions involve a dot product of log probabilities and assigned weights, but alternative formulations could provide more finely calibrated estimations. Additionally, the existing functions may not adequately capture complex dependencies between tokens, such as grammatical and semantic interactions De Marneffe and Nivre (2019). While MARS attempts to address this by weighting phrases rather than individual tokens, it only partially solves the problem and might fail to capture deeper dependencies. Consider the question, "What is the tallest building in the world?" and the model’s response: "The tallest building in the world might be Burj Khalifa with its lovely sight." Here, although the tokens "might" and "Burj Khalifa" may have high probabilities, "might" conveys uncertainty, suggesting that the model is uncertain despite the high probability of those tokens. An effective scoring function should recognize the interaction between "might" and "Burj Khalifa" and adjust the uncertainty accordingly. Additionally, the phrase "with its lovely sight" adds subjective opinion rather than factual reliability, yet it affects the overall meaning. Ignoring the probabilities of such tokens could improve the performance of the scoring function. Such important nuances are ignored by previous works. Lastly, both MARS and TokenSAR apply normalization on their weights $w(\mathbf{s},\mathbf{x},L,l)$, through methods like sum-normalization (TokenSAR) or softmax (MARS). Such design choices directly impact the UE output, potentially making the UE method converge to sub-optimal points.
Figure 2: Average accuracy and probability assignments of Llama2-7b-chat for specific entities in TriviaQA.
@@
Existing scoring functions directly utilize token probabilities, which may be biased against certain entity types Gallegos et al. (2024). To explore this issue, we conducted an experiment with Llama2-7b-chat Touvron et al. (2023) using TriviaQA Joshi et al. (2017). We posed TriviaQA questions to the model and analyzed the probabilities assigned to answer tokens representing different entity types like person names, organizations, and dates. Additionally, we assessed the model’s accuracy on questions whose ground truth answers are in these categories. As shown in Figure 2, though the model shows comparable accuracy for date and person entities, it assigns higher probabilities to date tokens. This finding suggests a positive bias towards date entities. We observed similar trends across other entity types. These differences in probability assignments highlight the need for recalibration across entities, which current scoring functions lack.
Performance in Different Languages.
MARS and TokenSAR rely on existing NLP tools for implementation. Specifically, TokenSAR uses a sentence similarity model Duan et al. (2024), and MARS relies on a QA evaluator model Bulian et al. (2022). These models may not be readily available for some low-resource languages. Moreover, the design of MARS and TokenSAR is primarily oriented towards English. This orientation may be challenging applied to languages that are morphologically distinct from English.
In the next section, we introduce a trainable scoring function, addressing these shortcomings.
@@ 4 LARS: Learnable Response Scoring
Intuition.
We develop a new scoring function that accounts for the semantic contributions of tokens in relation to the query, grasps biased probabilities, recognizes dependencies between tokens, and identifies other factors that may not be immediately apparent but are crucial for UE. Since manually designing a scoring function that has all these sophisticated properties would be extremely challenging as discussed in Section 3, we instead train a neural network with a transformer architecture that is capable of learning these properties directly from the data. An overview of the proposed approach is visualized in Figure 1.
Training Strategy.
@@
Let $f$ denote the scoring function, which accepts three arguments: the input prompt $\mathbf{x}=(x_{1},x_{2},\ldots,x_{N})$, the generated sequence $\mathbf{s}=(s_{1},s_{2},\ldots,s_{L})$, and the corresponding probability vector $\mathbf{p}=(p_{1},p_{2},\ldots,p_{L})$, where $p_{i}$ represents the probability of token $s_{i}$. The function $f$ outputs a real number $o$. This mapping captures crucial information: the meaning of the generated tokens ($\mathbf{s}$), their relevance to the context provided by the input prompt ($\mathbf{x}$), and the model’s confidence in each token via the probabilities ($\mathbf{p}$). In token probability-based methods, it is desirable for $o$ to be lower when the generation $\mathbf{s}$ is more likely to be incorrect, improving the model’s uncertainty estimation. To achieve this desired calibration, we make $f$ directly learnable through supervised data.
@@
We construct a calibration set to train our scoring function, $f_{w}$, which is parameterized by $w$. This calibration set comprises 4-tuples: input prompt $\mathbf{x}$, generated sequence $\mathbf{s}$, probability vector $\mathbf{p}$, and binary ground truth label $g$. The label $g$ indicates whether $\mathbf{s}$ is a correct response to $\mathbf{x}$. To optimize the parameters of $f_{w}$, we employ the binary cross-entropy loss, denoted by $L$, applied as follows: $L(f_{w}(\mathbf{x},\mathbf{s},\mathbf{p}),g).$ To train the scoring function $f_{w}$, we start with the pre-trained RoBERTa-base model Liu et al. (2019) and augment it by adding a linear layer that outputs a single logit.
Input Mapping.
Inputting text sequences $x$ and $s$ into a transformer model is straightforward, as we can leverage the standard text encoding strategy Vaswani (2017). However, encoding the probability information, which is a single real number for each token, poses a challenge due to its low dimensionality compared to the high-dimensional space of the model. To address this, we propose a novel input encoding strategy inspired by the class conditioning approach in conditional image generation van den Oord et al. (2016). We encode probability information to high-dimensional vectors by few-hot encoding. More specifically, we partition the probability range [0,1] to $k$ partitions. These partitions are mutually exclusive, cover the entire probability range, and are determined based on the quantiles of the probabilities in the calibration dataset. Given that the transformer model has an input dimension $d$, if $p_{i}$ falls in the range of $r$-th partition, we set its vector positions between $(r-1)\times\frac{d}{k}$ and $r\times\frac{d}{k}$ to 1, while all other positions are set to 0 (Figure 1 Mid Right). To ensure consistency with the model’s token embedding norms, we scale probability vectors by a fixed divisor and get the probability vector $\tilde{p_{i}}$. With this encoding strategy, we represent distinct probability ranges orthogonal to each other in high dimension. The input format of the LARS model is structured as follows (and visualized in Figure 1 Mid Left): initial prompt $\mathbf{x}$, followed by a series of response tokens $\mathbf{s}=(s_{1},s_{2},\ldots,s_{L})$. Each response token $s_{i}$ is immediately succeeded by its probability vector $\tilde{p_{i}}$.
UE Method Scoring Llama2-7b Llama3-8b Mistral-7b Gemma-7b Llama2-13b
Function AUROC PRR AUROC PRR AUROC PRR AUROC PRR AUROC PRR
TriviaQA Lex. Sim. - 0.647 0.374 0.683 0.483 0.720 0.517 0.597 0.227 0.611 0.314
# Sem. Gr. - 0.792 0.571 0.819 0.671 0.757 0.521 0.744 0.454 0.776 0.557
$\mathbf{p}$(True) - 0.616 0.267 0.842 0.733 0.805 0.653 0.517 0.023 0.650 0.392
SAPLMA - 0.741 0.484 0.736 0.541 0.785 0.614 0.658 0.373 0.757 0.594
Eccentricity - 0.812 0.629 0.853 0.756 0.818 0.664 0.764 0.496 0.813 0.633
Degree Matrix - 0.812 0.620 0.851 0.746 0.820 0.658 0.766 0.511 0.817 0.646
Confidence LNS 0.697 0.481 0.748 0.600 0.722 0.533 0.628 0.281 0.655 0.389
MARS 0.751 0.576 0.799 0.676 0.745 0.593 0.638 0.305 0.641 0.381
TokenSAR 0.747 0.572 0.792 0.674 0.747 0.584 0.688 0.386 0.728 0.527
LARS 0.851 0.760 0.872 0.817 0.844 0.759 0.819 0.690 0.846 0.766
SE LNS 0.795 0.627 0.835 0.733 0.810 0.670 0.749 0.475 0.800 0.617
MARS 0.797 0.645 0.835 0.742 0.810 0.681 0.749 0.482 0.794 0.615
TokenSAR 0.796 0.640 0.839 0.747 0.813 0.681 0.753 0.493 0.806 0.639
LARS 0.849 0.745 0.866 0.811 0.854 0.782 0.821 0.699 0.866 0.797
NaturalQA Lex. Sim. - 0.600 0.263 0.651 0.373 0.637 0.340 0.585 0.163 0.604 0.261
# Sem. Gr. - 0.705 0.379 0.736 0.448 0.675 0.283 0.686 0.276 0.709 0.377
$\mathbf{p}$(True) - 0.561 0.90 0.761 0.561 0.727 0.509 0.647 0.247 0.562 0.131
SAPLMA - 0.691 0.397 0.713 0.443 0.723 0.458 0.657 0.289 0.594 0.410
Eccentricity - 0.727 0.431 0.775 0.567 0.727 0.480 0.713 0.368 0.741 0.482
Degree Matrix - 0.727 0.435 0.771 0.554 0.732 0.483 0.715 0.358 0.742 0.487
Confidence LNS 0.677 0.384 0.697 0.449 0.666 0.390 0.610 0.189 0.648 0.338
MARS 0.699 0.411 0.717 0.477 0.678 0.407 0.615 0.198 0.631 0.311
TokenSAR 0.703 0.431 0.717 0.476 0.682 0.426 0.643 0.249 0.677 0.393
LARS 0.780 0.581 0.812 0.654 0.782 0.599 0.794 0.541 0.772 0.574
SE LNS 0.721 0.432 0.759 0.548 0.727 0.499 0.700 0.332 0.733 0.471
MARS 0.720 0.440 0.750 0.546 0.725 0.493 0.705 0.336 0.723 0.440
TokenSAR 0.721 0.443 0.756 0.544 0.726 0.498 0.700 0.340 0.736 0.485
LARS 0.772 0.569 0.794 0.638 0.778 0.591 0.785 0.548 0.779 0.583
Table 1: AUROC and PRR scores of UE methods on TriviaQA, NaturalQA.